# Boundary of the Moduli Space of Stable Cubic Fivefolds

### v6

@@ 6 Adjacency relations among strictly semistable components
The closed strata $\Phi_{1}$ and $\Phi_{7}$ are both zero-dimensional and adjacent. For $r\in\mathbb{Z}^{7}_{(0)}$, the slice $I(r)_{\geq 0}\cap I(r_{1})_{=0}$ is inclusion-maximal in $I(r_{1})_{=0}$ if and only if $r=r_{7}$, and $I(r)_{\geq 0}\cap I(r_{7})_{=0}$ is inclusion-maximal in $I(r_{7})_{=0}$ if and only if $r=r_{1}$. Moreover, for suitable specializations $\widetilde{\varphi}_{1},\widetilde{\varphi}_{7}$ (with all nonzero coefficients normalized to $1$), one has
$\lim_{t\to 0}\lambda_{7}(t)\cdot\widetilde{\varphi}_{1}$ $=\lim_{t\to 0}\lambda_{1}(t)\cdot\widetilde{\varphi}_{7}=x_{0}x_{5}^{2}+x_{1}^{2}x_{6}.$
Proposition 6.4.
The closed strata $\Phi_{2}$ and $\Phi_{6}$ are both zero-dimensional and adjacent. For $r\in\mathbb{Z}^{7}_{(0)}$, the slice $I(r)_{\geq 0}\cap I(r_{2})_{=0}$ is inclusion-maximal in $I(r_{2})_{=0}$ if and only if $r=r_{6}$, and $I(r)_{\geq 0}\cap I(r_{6})_{=0}$ is inclusion-maximal in $I(r_{6})_{=0}$ if and only if $r=r_{2}$. Moreover, for suitable specializations $\widetilde{\varphi}_{2},\widetilde{\varphi}_{6}$ (with all nonzero coefficients normalized to $1$), one has
$\lim_{t\to 0}\lambda_{6}(t)\cdot\widetilde{\varphi}_{2}$ $=\lim_{t\to 0}\lambda_{2}(t)\cdot\widetilde{\varphi}_{6}=x_{1}x_{3}x_{5}+x_{0}x_{3}x_{6}.$
Proposition 6.5.
The closed strata $\Phi_{3}$ and $\Phi_{12}$ are both zero-dimensional and adjacent. For $r\in\mathbb{Z}^{7}_{(0)}$, the slice $I(r)_{\geq 0}\cap I(r_{3})_{=0}$ is inclusion-maximal in $I(r_{3})_{=0}$ if and only if $r=r_{12}$, and $I(r)_{\geq 0}\cap I(r_{12})_{=0}$ is inclusion-maximal in $I(r_{12})_{=0}$ if and only if $r=r_{3}$. Moreover, for suitable specializations $\widetilde{\varphi}_{3},\widetilde{\varphi}_{12}$ (with all nonzero coefficients normalized to $1$), one has
$\lim_{t\to 0}\lambda_{12}(t)\cdot\widetilde{\varphi}_{3}$ $=\lim_{t\to 0}\lambda_{3}(t)\cdot\widetilde{\varphi}_{12}=x_{1}x_{4}^{2}+x_{2}^{2}x_{5}+x_{0}x_{3}x_{6}.$
Proposition 6.6.
The closed strata $\Phi_{8}$ and $\Phi_{19}$ are both two-dimensional and adjacent. For $r\in\mathbb{Z}^{7}_{(0)}$, the slice $I(r)_{\geq 0}\cap I(r_{8})_{=0}$ is inclusion-maximal in $I(r_{8})_{=0}$ if and only if $r=r_{19}$, and $I(r)_{\geq 0}\cap I(r_{19})_{=0}$ is inclusion-maximal in $I(r_{19})_{=0}$ if and only if $r=r_{8}$. Moreover, for suitable specializations $\widetilde{\varphi}_{8},\widetilde{\varphi}_{19}$ (with all nonzero coefficients normalized to $1$), one has
$\lim_{t\to 0}\lambda_{19}(t)\cdot\widetilde{\varphi}_{8}$ $=\lim_{t\to 0}\lambda_{8}(t)\cdot\widetilde{\varphi}_{19}$
$=x_{3}^{3}+x_{0}x_{4}^{2}+x_{0}x_{4}x_{5}+x_{0}x_{5}^{2}+x_{1}^{2}x_{6}+x_{1}x_{2}x_{6}+x_{2}^{2}x_{6}.$
Proposition 6.7.
The closed strata $\Phi_{9}$ and $\Phi_{15}$ are both two-dimensional and adjacent. For $r\in\mathbb{Z}^{7}_{(0)}$, the slice $I(r)_{\geq 0}\cap I(r_{9})_{=0}$ is inclusion-maximal in $I(r_{9})_{=0}$ if and only if $r=r_{15}$, and $I(r)_{\geq 0}\cap I(r_{15})_{=0}$ is inclusion-maximal in $I(r_{15})_{=0}$ if and only if $r=r_{9}$. Moreover, for suitable specializations $\widetilde{\varphi}_{9},\widetilde{\varphi}_{15}$ (with all nonzero coefficients normalized to $1$), one has
$\lim_{t\to 0}\lambda_{15}(t)\cdot\widetilde{\varphi}_{9}$ $=\lim_{t\to 0}\lambda_{9}(t)\cdot\widetilde{\varphi}_{15}=x_{3}^{3}+x_{0}x_{4}x_{5}+x_{1}x_{2}x_{6}+x_{0}x_{3}x_{6}.$
Proposition 6.8.
The closed strata $\Phi_{10}$ and $\Phi_{17}$ are both two-dimensional and adjacent. For $r\in\mathbb{Z}^{7}_{(0)}$, the slice $I(r)_{\geq 0}\cap I(r_{10})_{=0}$ is inclusion-maximal in $I(r_{10})_{=0}$ if and only if $r=r_{17}$, and $I(r)_{\geq 0}\cap I(r_{17})_{=0}$ is inclusion-maximal in $I(r_{17})_{=0}$ if and only if $r=r_{10}$. Moreover, for suitable specializations $\widetilde{\varphi}_{10},\widetilde{\varphi}_{17}$ (with all nonzero coefficients normalized to $1$), one has
$\lim_{t\to 0}\lambda_{17}(t)\cdot\widetilde{\varphi}_{10}$ $=\lim_{t\to 0}\lambda_{10}(t)\cdot\widetilde{\varphi}_{17}=x_{3}^{3}+x_{1}x_{3}x_{5}+x_{0}x_{4}x_{5}+x_{1}x_{2}x_{6}.$
Proposition 6.9.
The closed strata $\Phi_{11}$ and $\Phi_{21}$ are both four-dimensional and adjacent. For $r\in\mathbb{Z}^{7}_{(0)}$, the slice $I(r)_{\geq 0}\cap I(r_{11})_{=0}$ is inclusion-maximal in $I(r_{11})_{=0}$ if and only if $r=r_{21}$, and $I(r)_{\geq 0}\cap I(r_{21})_{=0}$ is inclusion-maximal in $I(r_{21})_{=0}$ if and only if $r=r_{11}$. Moreover, for suitable specializations $\widetilde{\varphi}_{11},\widetilde{\varphi}_{21}$ (with all nonzero coefficients normalized to $1$), one has
$\lim_{t\to 0}\lambda_{21}(t)\cdot\widetilde{\varphi}_{11}$ $=\lim_{t\to 0}\lambda_{11}(t)\cdot\widetilde{\varphi}_{21}$
$=x_{2}^{3}+x_{2}^{2}x_{3}+x_{2}x_{3}^{2}+x_{3}^{3}+x_{2}^{2}x_{4}+x_{2}x_{3}x_{4}+x_{3}^{2}x_{4}+x_{2}x_{4}^{2}+x_{3}x_{4}^{2}+x_{4}^{3}.$
Proposition 6.10.
The closed strata $\Phi_{14}$ and $\Phi_{20}$ are both one-dimensional and adjacent. For $r\in\mathbb{Z}^{7}_{(0)}$, the slice $I(r)_{\geq 0}\cap I(r_{14})_{=0}$ is inclusion-maximal in $I(r_{14})_{=0}$ if and only if $r=r_{20}$, and $I(r)_{\geq 0}\cap I(r_{20})_{=0}$ is inclusion-maximal in $I(r_{20})_{=0}$ if and only if $r=r_{14}$. Moreover, for suitable specializations $\widetilde{\varphi}_{14},\widetilde{\varphi}_{20}$ (with all nonzero coefficients normalized to $1$), one has
$\lim_{t\to 0}\lambda_{20}(t)\cdot\widetilde{\varphi}_{14}$ $=\lim_{t\to 0}\lambda_{14}(t)\cdot\widetilde{\varphi}_{20}$
$=x_{0}x_{3}x_{5}+x_{1}x_{3}x_{5}+x_{0}x_{3}x_{6}+x_{1}x_{3}x_{6}.$
Collecting Propositions 6.36.10, we obtain Theorem 6.1.

### v7

@@ 8.5.1 Filter 6: spectra of isolated singularities (and the last three pairs)
The inclusion $\Phi_{5}\subset\Phi_{16}$ is impossible.
Proof.
Assume $\Phi_{5}\subset\Phi_{16}$. Then the closed orbit representative $X_{5}$ lies in the closure of $\Phi_{16}$. Choose a one-parameter degeneration $X_{t}\in\Phi_{16}$ with limit $X_{0}\cong X_{5}$. For $t\neq 0$, the variety $X_{t}$ has two isolated singular points of type $E_{20}$ (Proposition 5.23); the limit $X_{0}$ has two isolated singular points of type $E_{21}$. Working in disjoint analytic neighborhoods of the singular points, we obtain families of isolated hypersurface singularities and can apply Theorem 8.16. This contradicts the strict inequality $\bigl(\mathrm{Sp}(E_{21})\cap(c,c+1]\bigr)<\#\bigl(\mathrm{Sp}(E_{20})\cap(c,c+1]\bigr)$. ∎
(B) The remaining three pairs.
After Proposition 8.18, the only remaining candidate inclusions are
$\Phi_{6}\subset\Phi_{5},\qquad\Phi_{6}\subset\Phi_{16},\qquad\Phi_{9}\subset\Phi_{16}.$
We exclude these by comparing the generic connected stabilizer tori.
Lemma 8.19 (Generic stabilizer type obstruction).
Let $G$ act on a projective variety $X$ and let $S\subset X$ be a $G$-stable irreducible subset. Assume that a general point of $S$ has connected stabilizer a one-dimensional torus $T\cong\mathbb{C}^{*}$. If $S\subset S^{\prime}$ for another $G$-stable irreducible subset $S^{\prime}\subset X$ whose general point has connected stabilizer a one-dimensional torus $T^{\prime}$, then $T^{\prime}$ is conjugate to $T$ in $G$. In particular, if both $T$ and $T^{\prime}$ are diagonal one-parameter subgroups of $\mathrm{SL}_{7}$ then their weight multisets must agree.
Proof.
Under specialization the connected stabilizer can only increase. Hence for $S\subset S^{\prime}$ the generic torus stabilizer of $S^{\prime}$ embeds into the generic torus stabilizer of $S$. Since both are one-dimensional tori, the embedding is an isomorphism, i.e. the two tori are conjugate. For diagonal one-parameter subgroups of $\mathrm{SL}_{7}$, conjugacy is equivalent to equality of the weight multiset. ∎
In our setting, the closed orbit representatives $X_{k}$ are fixed by the diagonal one-parameter subgroups $\lambda_{k}$ listed in Section 2. For $k\in\{5,6,9,16\}$ these are
$\lambda_{5}(t)$ $=\mathrm{diag}(t^{4},t^{2},t^{1},t^{0},t^{-1},t^{-2},t^{-4}),$
$\lambda_{6}(t)$ $=\mathrm{diag}(t^{5},t^{3},t^{2},t^{1},t^{-1},t^{-4},t^{-6}),$
$\lambda_{9}(t)$ $=\mathrm{diag}(t^{2},t^{2},t^{0},t^{0},t^{-1},t^{-1},t^{-2}),$
$\lambda_{16}(t)$ $=\mathrm{diag}(t^{2},t^{1},t^{0},t^{0},t^{0},t^{-1},t^{-2}).$
Their weight multisets are pairwise distinct, so Lemma 8.19 rules out the three remaining inclusions.
Corollary 8.20.
After applying Filters 1–6, there are no remaining candidate inclusions $\Phi_{k}\subset\Phi_{\ell}$ with $k\neq\ell$. In particular, Theorem 3.5 holds.
9 Concluding Remarks and Future Directions
The results of this paper turn the strictly semistable GIT boundary for cubic fivefolds into an explicit and tractable set of test degenerations. Concretely, we obtain: the list of 2121 $\mathrm{SL}(7)$-inequivalent boundary families (Theorem 3.6), explicit polystable normal forms and component dimensions (Section 4 and Table 2), a complete description of the singular loci of the closed-orbit representatives (Section 5 and Table 3), the six isolated analytic types that occur on the boundary (Table 4), the sharp minimal-exponent equalities at the critical threshold (Theorems 6.1 and 6.4), and the sparse Kirwan wall-crossing adjacency graph among boundary components (Section 7 and Theorem 7.2). Together with the non-inclusion statement (Theorem 3.5), this provides a concrete “dictionary” for the boundary geometry of the moduli space.
We close by outlining several directions where such explicit boundary models should be useful. A long-term motivation is to leverage these degenerations in Hodge-theoretic and integrable-systems approaches to the intermediate Jacobians of cubic fivefolds (Section 9.3). To make that program viable, however, one needs a substantially refined understanding of (i) the deformation and moduli theory of the extremal isolated singularities (Section 9.1), (ii) the resulting limit mixed Hodge structures for boundary degenerations (Section 9.2), and (iii) the extent to which the present methodology scales to higher-dimensional cubic hypersurfaces (Section 9.4). We also record a conjectural structural principle suggested by our computations (Section 9.5).
9.1 Detailed analysis of extremal cubic fivefold singularities
Isolated singular points occur only for a proper subset of the boundary families (Table 3), and, on general closed-orbit representatives, every isolated boundary germ belongs to one of the six quasi-homogeneous analytic types listed in Table 4. A key outcome of Section 6 is that each of these isolated types realizes the equality case in Park’s threshold: Theorem 6.1 shows
$\widetilde{\alpha}_{x}(X)=\frac{7}{3}=\frac{n+1}{d}\qquad\text{for }(n,d)=(6,3)$
at every isolated boundary point. Motivated by this sharp borderline behavior, we singled out these six analytic types as extremal (Definition 6.3).
Beyond their role in the GIT classification, these singularities appear to be a natural and rigid collection of “boundary” local models for cubic fivefolds, and they invite a more intrinsic analysis. Several concrete problems seem especially compelling:
  • Local deformation theory and comparison with wall-crossing. A systematic study of the miniversal deformation spaces (semi-universal unfoldings) of the six extremal types should clarify the local analytic structure of the moduli space near boundary points. Since the singularities are quasi-homogeneous, one expects the Milnor and Tjurina numbers to coincide, and the deformation space can be approached via explicit Tjurina algebras. A natural goal is to understand how (semi)stability conditions cut out chambers inside these bases and how the resulting local pictures reflect Kirwan wall-crossing (Definition 7.1, Theorem 7.2).
  • Connections to $K$-stability and $K$-moduli. With the rapid development of $K$-moduli for Fano varieties, it is natural to ask how the classical GIT compactification compares to the $K$-moduli compactification for cubic fivefolds. While minimal exponents are Hodge-theoretic invariants and do not directly encode discrepancies, the uniform inequality $\widetilde{\alpha}_{x}(X)=\frac{7}{3}>1$ for the extremal isolated types suggests that these singularities should be relatively mild from several birational and Hodge-theoretic perspectives. It would therefore be very interesting to determine for each explicit extremal local model whether it is klt, and more broadly to test local $K$-stability by computing normalized volumes and related stability thresholds.
  • Finer Hodge-theoretic and topological invariants. Since these extremal germs govern the most degenerate isolated limits on the GIT boundary, their refined invariants (Hodge spectrum, monodromy, Milnor lattice, Brieskorn lattice, etc.) should feed directly into computations of limiting Hodge data for global degenerations (cf. Section 9.2).
9.2 Limit mixed Hodge structures and moduli compactifications
For cubic threefolds and fourfolds, a recurring theme is that explicit GIT compactifications interact in a meaningful way with Hodge-theoretic compactifications of period domains (e.g. [CG72, Laz09]). For cubic fivefolds, understanding the degeneration of periods and intermediate Jacobians [Col86] requires an analysis of the associated limit mixed Hodge structures (LMHS).
From this perspective, two kinds of input are indispensable: one needs explicit local models of singular fibers, and one needs control over how boundary strata meet. The present paper provides precisely this data in a concrete form: explicit closed-orbit boundary representatives (Table 2), their singular loci (Section 5, Table 3), and the adjacency relations among boundary components under Kirwan wall-crossing (Section 7, Theorem 7.2). Combined with the local invariants of the extremal isolated types (Section 9.1), this should make it feasible to construct transverse slices to boundary strata and to compute local Picard–Lefschetz transformations and monodromy cones in explicit families. A natural medium-term objective is therefore a systematic computation of LMHS for selected degenerations built from our $21$ boundary models, with an eye toward comparisons between the GIT boundary and Hodge-theoretic compactifications.
9.3 Towards the Schottky problem for cubic fivefolds via integrable systems
Let $X\subset\mathbb{P}^{6}$ be a smooth cubic fivefold. Its intermediate Jacobian $J(X)$ is a principally polarized abelian variety (PPAV) of dimension $21$, and a Torelli-type statement is available in this setting (see [Voi86]). Thus the period map may be viewed as embedding the $35$-dimensional moduli space of smooth cubic fivefolds into the $231$-dimensional PPAV moduli space $\mathcal{A}_{21}$. A basic and widely open problem is to characterize, or even effectively constrain, this $35$-dimensional locus inside $\mathcal{A}_{21}$.
In the classical Schottky problem for Jacobians of curves, integrable systems play a decisive role: Shiota’s theorem characterizes Jacobians among PPAVs via the KP hierarchy [Shiota86]. This motivates the question of whether the locus of intermediate Jacobians of cubic fivefolds admits an analogous characterization in terms of modular/differential constraints on theta functions. At present, even formulating a plausible candidate hierarchy is nontrivial, and any integrable-systems approach must confront degenerations.
Here the boundary classification becomes relevant. Theta-functional constraints are typically most informative when one understands their behavior under degeneration, where period matrices and theta functions develop controlled asymptotics. Our $21$ explicit boundary families (Theorem 3.6 and Table 2), together with the specialization relations encoded by Theorem 7.2, provide a concrete set of test degenerations and a combinatorial skeleton for how boundary strata meet. A natural open-ended program is therefore: use these explicit degenerations to compute LMHS and limiting period data (as in Section 9.2), and then investigate whether the resulting locus in $\mathcal{A}_{21}$ satisfies distinctive modular or differential constraints of an integrable-systems flavor.
9.4 Extremal singularities on cubic sixfolds and higher dimensions
The computational and structural pipeline used in this paper (maximal strictly semistable supports via convex geometry, Luna’s centralizer reduction, and polystability checks via the convex-hull criterion or the Casimiro–Florentino criterion) relies on general features of reductive group actions and is therefore, in principle, portable to higher-dimensional cubic hypersurfaces, albeit with rapidly growing combinatorial complexity.
An immediate target is the GIT compactification for cubic sixfolds ($X\subset\mathbb{P}^{7}$). Park’s stability criterion [Par25, Theorem A] identifies the critical minimal exponent threshold as $\frac{n+1}{d}$ for degree $d$ hypersurfaces in $\mathbb{P}^{n}$; in particular, for cubic sixfolds $(n,d)=(7,3)$ the critical value is $\frac{8}{3}$. Our results for cubic fivefolds show that the isolated boundary singularities realize the equality case $\widetilde{\alpha}_{x}(X)=\frac{7}{3}$ (Theorem 6.1), suggesting a broader principle: in higher dimensions, the strictly semistable boundary may again be governed by a distinguished collection of quasi-homogeneous isolated singularities sitting exactly at the critical threshold. Identifying such “extremal” types for cubic sixfolds would be a natural next step toward a higher-dimensional boundary dictionary.
9.5 A hypersurface version of maximal polystability
One striking uniform phenomenon emerged throughout our explicit classification: for every maximal $T$-strictly semistable support $I(\mathbf{r}_{k})_{\geq 0}$ produced by the convex-geometric algorithm, the associated $1$-PS degeneration
$\phi_{k}\;=\;\lim_{t\to 0}\lambda_{\mathbf{r}_{k}}(t)\cdot f_{k}$
of a general polynomial $f_{k}$ with $\mathrm{Supp}(f_{k})=I(\mathbf{r}_{k})_{\geq 0}$ landed on a polystable point for the residual action of the corresponding centralizer. In other words, once the support is maximal among strictly semistable ones, the weight-0 initial form produced by the limiting process appears to have closed orbit automatically.
The fact that this automatic polystability holds flawlessly without a single exception across all 21 highly diverse boundary families computed in Section 4—despite the wide variation in their residual centralizers, ranging from tori to complicated non-abelian groups—provides overwhelming empirical evidence.
Heuristically, the convex-geometric maximality of the support dictates a combinatorial rigidity: one cannot add further monomials without making the support unstable. Intuitively, this extremality forces the barycenter to be strictly ”trapped” in the relative interior of the weight polytope for the residual action, which exactly translates to polystability.
Motivated by this evidence, we record the following hypersurface-level formulation, tailored to the natural $\mathrm{SL}(n{+}1)$-action on spaces of homogeneous polynomials.
Conjecture 9.1 (Hypersurface Maximal Polystability).
Fix integers $n\geq 1$ and $d\geq 2$, and let
$G=\mathrm{SL}(n{+}1),\qquad W=\mathrm{Sym}^{d}(\mathbb{C}^{n{+}1})\cong\mathbb{C}[x_{0},\dots,x_{n}]_{d}$
with the natural linear $G$-action. Fix the diagonal maximal torus $T\subset G$. Let
$I=\mathbb{Z}^{n+1}_{(d)}\cap\mathbb{Z}^{n+1}_{\geq 0},\qquad\eta=\Bigl(\tfrac{d}{n+1},\dots,\tfrac{d}{n+1}\Bigr)$
be the exponent simplex and its barycenter. For $\mathbf{r}\in\mathbb{Z}^{n+1}_{(0)}$ define
$I(\mathbf{r})_{\geq 0}=\{\mathbf{i}\in I\mid\mathbf{r}\cdot\mathbf{i}\geq 0\},\qquad I(\mathbf{r})_{=0}=\{\mathbf{i}\in I\mid\mathbf{r}\cdot\mathbf{i}=0\}.$
We say that $I(\mathbf{r})_{\geq 0}$ is a maximal $T$-strictly semistable support if $\eta\in\mathrm{Conv}\bigl(I(\mathbf{r})_{\geq 0}\bigr)$, $\eta\in\partial\mathrm{Conv}\bigl(I(\mathbf{r})_{\geq 0}\bigr)$ (relative boundary inside the affine hyperplane $\mathrm{wt}=d$), and $I(\mathbf{r})_{\geq 0}$ is maximal for these properties with respect to inclusion among subsets of $I$ of the form $I(\mathbf{r}^{\prime})_{\geq 0}$. Let $\lambda_{\mathbf{r}}:\mathbb{G}_{m}\to T$ be the $1$-PS given by $\lambda_{\mathbf{r}}(t)=\mathrm{diag}(t^{r_{0}},\dots,t^{r_{n}})$. For a general form
$f=\sum_{\mathbf{i}\in I}a_{\mathbf{i}}x^{\mathbf{i}}\in W\quad\text{with}\quad\mathrm{Supp}(f)=I(\mathbf{r})_{\geq 0},$
consider the $1$-PS limit
$f_{0}\;:=\;\lim_{t\to 0}\lambda_{\mathbf{r}}(t)\cdot f\;=\;\sum_{\mathbf{i}\in I(\mathbf{r})_{=0}}a_{\mathbf{i}}x^{\mathbf{i}}\;\in\;W^{\lambda_{\mathbf{r}}(\mathbb{G}_{m})}.$
Then $f_{0}$ is polystable.
Possible strategy for a proof. One may hope to prove Conjecture 9.1 by combining Luna’s centralizer reduction [Lun75] with the Casimiro–Florentino criterion [CF12] (or, in toric cases, the convex-hull criterion, Theorem 4.1). Luna reduction translates closedness of the $G$-orbit of the $1$-PS limit $f_{0}$ into polystability for the residual action of the centralizer $C_{G}(\lambda(\mathbb{G}_{m}))$ on the fixed locus $W^{\lambda(\mathbb{G}_{m})}$. After choosing a maximal torus in the centralizer, the Casimiro–Florentino/convex-hull criteria reduce polystability to the existence of a strictly positive “balancing” linear identity in the real character space.
In the hypersurface case $W=\mathrm{Sym}^{d}\mathbb{C}^{n+1}$, this can be phrased as the existence of positive coefficients $a_{\alpha}>0$ (summing to $1$) such that
$\sum_{\alpha\in\mathrm{Supp}(f_{0})}a_{\alpha}\,\alpha\;=\;\frac{d}{n+1}(1,\ldots,1).$
The conjecture asserts that the maximal-support hypothesis forces such a strictly positive solution, equivalently that the relevant weight polytope already contains the barycenter in its interior for the residual action. If true, this would streamline the search for closed orbits in many hypersurface GIT problems: once maximal strictly semistable supports are enumerated, the residual polystability check would become essentially automatic.